Richie Richs wealth is $500,000. He makes $40,000 a year, on which he is taxed 30%. Richies weath ini des some o uprent, which depreciates by $5,000 a year Based on this which of the olowing statements is correct? Richies maximum amount of consumption expendture possible is $23,000 Richies net income is $28,000 团Richie decides to spend 60% of his net income on consumption and the rest on nvestment, then his invesment-S9200. Richies disposable income is $40,000.
Answer #1

Richie's net income is $28,000.

Net income = Income earned after taxes = 40,000 - 30%*(40,000) = 40,000 - 12,000 = 28,000
So, Richie's net income is $28,000.